# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v4

@@ 8.3 Sim-to-Real Gap Solutions
The sim-to-real problem is a widespread issue across machine learning, not limited to manipulation (Zhao et al. (2020)). The goal is to successfully transfer the policy from the simulation (source domain) to the real world (target domain). The gap in the manipulation tasks between the simulation and the real-world includes two main types: visual gap and dynamic gap. Visual gap refers to the difference between the vision information produced by the renderer and the vision information in the real world. The dynamic gap consists of several factors. First, there is a difference between the physics engine used in simulations and real-world physics. Second, the properties of objects, including robots, contribute to the object dynamic gap. Lastly, there is a control gap in robots, such as variations in static errors caused by different PID parameters. Currently, there are three main approaches to address sim-to-real gap: system identification, domain randomization, and transfer learning (Zhao et al. (2020)).
System identification (Kristinsson and Dumont (1992)) aims to create an accurate mathematical model for a physical system to make the simulator more realistic. However, it is impossible to accurately build models of complex environments in simulators.
@@ 8.4 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (Mandlekar et al. (2023)) and DexMimicGen (Jiang et al. (2024c)) change the positions and orientations of objects, while CACTI (Mandi et al. (2022)) adds new, artificial objects to the scene. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@
Following Fig. 8, LLMs can generate credible descriptions or code for task scenes. VGMs produce 3D object meshes and render textures. Nonetheless, the validity of the generated task scenes must be ultimately assessed by VLMs.
@@ 9.1 What is the framework for general manipulation?
9.2 How to design post-conditions detection and post-hoc correction?
The current data collection only focuses on gathering successful task execution data, ignoring the collection of data related to failed task executions. However, if data on failed task executions are collected and annotated with corresponding error reasons, it would be possible to train a model to both determine task execution success and analyze the reasons for task execution failure. For instance, RoboVQA (Sermanet et al. (2023)) utilizes crowdsourced data collection and trains videoCoCo (Yan et al. (2022)), employing a video-language model to analyze task execution description. Similarly, a post-conditions detection model could utilize a video-language model for training, taking manipulation videos as input to output whether tasks are successfully executed and to analyze the reasons for task execution failure. Post-hoc correction could then generate corrective action sequences based on the reasons for task execution failure and the task objectives, which would be handed over to a policy to generate corresponding corrective actions.
@@ 9.3 What kind of learning capability should a general manipulation framework possess?
9.4 The strengths and weaknesses of current RFMs.
This paper categorizes foundation models trained using manipulation datasets as RFMs. RFMs significantly enhance the policy’s generalization. For instance, LEO (Huang et al. (2023a)) directly outputs target pose, while RoboFlamingo (Li et al. (2023d)) directly outputs delta pose. Directly outputting target pose is easier to train compared to delta pose, although delta pose better aligns with human behavior. Currently, RFMs are fine-tuned using pre-trained models, but this deprives them of self-exploratory learning. Designing an architecture for RFMs is crucial. Utilizing foundation models to aid reinforcement learning still holds many unexplored areas, such as using foundation models to assist robots in reinforcement learning within real-world environments.
Training RFMs simultaneously requires a large amount of training data. Current data collection methods include gathering data in simulation, using skill library to generate data, teleoperation, and imitation learning from human videos. While collecting data in simulation is cost-effective, it suffers from the sim-to-real gap. In order to generate data using skill library, it is necessary to construct an initial skill library. Teleoperation can address the sim-to-real gap but comes with high hardware costs. Imitation learning from human videos involves mapping human trajectories from videos to robot space, leading to issues with retargeting. Which of these methods can effectively lead to general manipulation remains to be explored.
@@ 9.5 How to use internet-scale video data for RFMs?
Compared to single-frame images and language data on the internet, internet videos contain information on the physics and dynamics of the world, as well as on human behaviors and actions (Chandrasegaran et al. (2024)). This information is precisely what is required for manipulation tasks. Therefore, in this subsection, we introduce the use of internet-scale video data for training robotic foundation models.
9.5.1 Video Dataset.
Methods to obtain language annotations are divided into manual and automated captions. Manual captions are created by humans labeling video content. Automated captions include four types: (1) Automatic Speech Recognition (ASR), which converts audio in videos to text (Xue et al. (2022)). (2) Alt-text, which collects captions from HTML alt-text attributes in web images and videos, like descriptions, tags, and titles (Bain et al. (2021)). (3) Transfer, which starts with a set of image-caption pairs. Then, captions are matched to video clips with similar frames (Nagrani et al. (2022)). (4) Foundation Models, which use pre-trained models to get captions. For example, VLMs provide single-frame image captions, while LLMs filter out inconsistent captions across frames (Blattmann et al. (2023)). Owing to recent advancements in language annotation techniques, most widely used internet video datasets incorporate language annotations, such as InternVid (Wang et al. (2023d)), HD-VILA-100M (Xue et al. (2022)), YT-Temporal-180M (Zellers et al. (2021)), WTS-70M (Shvetsova et al. (2025)), HowTo100M (Miech et al. (2019)), WebVid-10M (Nan et al. (2024)), and VideoCC3M (Yan et al. (2022)).
9.5.2 Video Dataset for Robotic Foundation Models.
@@
Transferring human actions from video to robot control is similar to cross-embodiment learning (Zakka et al. (2022)). Three key questions require focus: what information from video dataset can be used, how to extract these useful information and how to utilize the extracted information to enhance or train robotic foundation models. Regarding the first question, there are six main types of information to convert from video datasets: (1) Pose, such as capturing human hand poses and retargeting them to dexterous hand poses (Shaw et al. (2023); Qin et al. (2022)). (2) Affordance, including grasp locations on objects and post-grasp waypoints (Mendonca et al. (2023)). (3) Motion information, explicitly includes keypoints trajectories of objects and human hand during actions (Xiong et al. (2021); Yuan et al. (2024); Wen et al. (2023b)) and implicitly includes using VQ-VAE (Van Den Oord et al. (2017)) to generate a codebook for latent delta action (Ye et al. (2024)). (4) Environment transition dynamic information, such as predicting hindsight images after completing the current action (Wu et al. (2023a); Cheang et al. (2024); Yang et al. (2023b)). (5) Semantic information, such as descriptions of current task steps (Wang et al. (2024a)) and task instruction (Jain et al. (2024)). (6) Spatial and texture information, such as MVP (Radosavovic et al. (2023)) suggests using masked autoencoding (He et al. (2022)) for improving visual reconstruction.
@@
Regarding the second question, the methods of language annotation have been introduced above. At the same time, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023b)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)). When labeling pose for videos, research on human pose detection and retargeting becomes crucial (Shaw et al. (2023); Qin et al. (2022); Li et al. (2024)). When adding various labels to the video dataset, different training objectives can be used to extract features from the video dataset, such as MAE (Radosavovic et al. (2023)), contrastive learning (Ma et al. (2022)), time contrastive learning (Ma et al. (2023a)), temporal-difference learning (Bhateja et al. (2023)), video prediction objective(Du et al. (2024)), affordance prediction objective(Mendonca et al. (2023)), video-language alignment objective (Nair et al. (2022)), action motion objective(Yuan et al. (2024)) or combinations of these objectives (Karamcheti et al. (2023); Zhou et al. (2021)).
@@
The current robotic foundation models primarily use two learning methods: imitation learning and reinforcement learning. Therefore, the discussion on the third issue focuses on leveraging prior knowledge from video datasets in these two methods. As for imitation learning, when the robotic foundation model outputs pose and the video dataset annoated with pose label, the video dataset can be directly used as training data for the robotic foundation model (Shaw et al. (2023); Qin et al. (2022)). When leveraging affordance information, motion information, environment transition dynamics information, semantic information, spatial and texture information, it is essential to employ GMM & CEM (Mendonca et al. (2023)), Inverse Dynamic Model (IDM) (Du et al. (2024); Ye et al. (2024); Wen et al. (2023b)), and Decoder (Wang et al. (2023a); Xiao et al. (2022b); Cheang et al. (2024); Wu et al. (2023a)) to transform these information into actions. Compared to other types of information, using semantic information treats the video as task instruction rather than observation (Jain et al. (2024); Shah et al. (2023); Jang et al. (2022)). At the same time, semantic information can also be used to organize tasks into a hierarchy of skills (Wang et al. (2024a)).
@@
9.6 How to design low-cost hardware for collecting large-scale real-world manipulation data?
Demonstration data plays a signification role in robotic manipulation, especially for imitation learning. A natural approach to gather such demonstrations is human teleoperation. Recently, there are increase number of excellent works in low-cost teleoperation hardware that can collect high standard demonstration data.
9.6.1 Teleoperation Types.
9.6.2 Human Motion Measurement and Visual Feedback.
Above systems work in cartesian space, which needs inverse kinematic (IK) solver and off-the-shelf IK often suffering from fails when operating near singularities of the robot. Although some bilateral teleoperation systems use haptic feedback to provide a tangible sense of the robot’s kinematic constraints, they do not address the challenges of very tight operational spaces (Silva et al. (2009)). Therefore, multi-joint encoder teleoperation devices can solve the IK problem by working in the joint space. The current design of multi-joint teleoperation devices is mainly divided into isomorphic and non-isomorphic devices (Wu et al. (2023c)). Isomorphic devices refer to teleoperation systems using standard servo-based robotic arms to control manipulators with similar size and kinematics (Zhao et al. (2023a)), while non-isomorphic devices use such arms to control manipulators with different size and kinematic properties. Non-isomorphic devices use kinematically equivalent structures based on DH parameters to map joint spaces between different properties (Wu et al. (2023c)). Therefore, non-isomorphic devices can enable a single teleoperation device to operate robots with different configurations. AirExo (Fang et al. (2024)) expands this low-cost and scalable platform into a wearable device to collect cheap in-the-wild demonstrations at scale.
9.6.3 Summary.
9.7 How to collect large-scale diversity and realism manipulation data in simulation?
In the simulation, generating manipulation data involves three steps. First, create the simulation scene. Second, collect demonstration data based on the task. Third, apply data augmentation to the collected data. In Section 8, we introduce the simulator, scene and demonstration generation, sim-to-real gap solutions, and dataset augmentation. However, achieving diversity and realism in simulation manipulation data remains a significant challenge. Diversity is mainly reflected in scene diversity, task diversity and data augmentation. Realism focuses on the reality of simulations.
As described in subsection 8.2, current scene reconstruction approaches are mainly divided into Real-to-Sim methods and Automated Generation methods. Compared to Real-to-Sim methods, Automated Generation methods are more effective for producing large-scale diverse scenes (Deitke et al. (2022)). Therefore, this discussion focuses on the impact of Automated Generation methods on scene diversity.
Scene diversity primarily includes the diversity of scene layouts, such as floor plans and object placements, as well as the diversity of objects (Deitke et al. (2022)). Regarding the automatic generation of floor plans, current approaches are either rule-based (Deitke et al. (2022)) or learning-based (Yang et al. (2024b)). However, they do not leverage the commonsense knowledge embedded in foundation models. Such knowledge can greatly enhance the naturalness of room layouts. For example, bedrooms are often connected to adjacent bathrooms through a door. Similarly, this commonsense knowledge applies to object layouts as well. For instance, knives are typically placed on the right side of plates, while forks are on the left (Ding et al. (2023)). Regarding the diversity of objects, although there are many existing 3D object assets (Chang et al. (2015); Deitke et al. (2023); Li et al. (2023a); Geng et al. (2023a); Xiang et al. (2020); Liu et al. (2022b); Calli et al. (2017)), their quantity is far from sufficient to cover the variety of real-world objects. As a result, generative models have gained popularity as a method for generating 3D objects. However, the performance of generative models is also limited by the shortage of current 3D object assets cannot cover objects in the real world. To address this issue, data cleaning techniques or manual supervision are needed to filter and select high-quality generated object assets. At the same time, generative models mainly focus on rigid and articulated objects and research on deformable objects remains insufficient (Sundaresan et al. (2022)).
Table 2: Representative Low-cost Hardware Works.
Teleoperation Device Teleoperation Type Human Motion Measurement Feedback Cost
Aloha(Zhao et al. (2023a)) Online Teleoperation Joint Copy Third-person View \$2000(include robot)
GELLO(Wu et al. (2023c)) Online Teleoperation Joint copy Third-person View \$300
Human Plus(Fu et al. (2024)) Online Teleoperation End-effectors mapping Third-person View \$30(A RGB Camera)
Transteleop(Li et al. (2020)) Online Teleoperation End-effectors mapping Third-person View \$170
OPEN TEACH(Iyer et al. (2024)) Online Teleoperation End-effectors mapping First-person View \$500
Open-TeleVision(Cheng et al. (2024)) Online Teleoperation End-effectors mapping First-person View \$3499
AirExo (Fang et al. (2024)) Offline Teleoperation Joint copy - \$600
UMI(Chi et al. (2024)) Offline Teleoperation End-effectors mapping - \$371
DexCap(Chi et al. (2024)) Offline Teleoperation End-effectors mapping - \$4000
VideoDex(Chi et al. (2024)) Offline Teleoperation End-effectors mapping - -
As for task diversity, many studies currently utilize LLMs to generate instructions, followed by the use of skill libraries to create demonstrations (Wang et al. (2023c, e)). However, generating diverse task instructions using foundation models presents challenges in ensuring rationality. For example, integrating observations from specific scenarios to identify existing objects and their affordances is crucial. This approach helps make the instructions more feasible and actionable.
Data augmentation plays a crucial role in enhancing data diversity. Current methods are mainly divided into scene-level and object-level approaches. However, the reliability of data augmentation still needs validation. For example, MimicGen (Mandlekar et al. (2023)) filters data generation attempts based on task success. For scene-level scenarios, dynamic changes should be incorporated, such as variations in the background. Additionally, obstacle avoidance considerations should also be added for scene-level scenarios.
As for the reality of simulation, it mainly includes rendering realism and physical realism. In terms of rendering realism, BEHAVIOR-1K (Li et al. (2023a)) highlights that Omniverse offers the highest rendering performance. For physical realism, current simulations require physical parameter identification to ensure objects in the simulation have similar physical properties to real-world objects. The primary methods for physical parameter identification include estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)), and estimation from observations using foundation models (Gao et al. (2023)). Among these, estimation from demonstrations appears more effective. Demonstrations inherently contain interaction information and can also assist policy training. However, improving the hardware performance for collecting demonstrations remains essential.
9.8 Comparative analysis of 2D and 3D-based methods.
Current manipulation tasks are mainly divided into high-level and low-level. High-level tasks involve decision-making, such as the hierarchy of skills. Low-level tasks focus on execution, like policies. For high-level manipulation task, 3D observation has stronger spatial reasoning capabilities compared to 2D observation (Chen et al. (2024b)). It can recognize quantitative relationships of physical objects, such as distances or size differences. For low-level manipulation task, DP3 has shown that diffusion policies with 3D input achieve higher success rates compared to 2D image and the point cloud format performs best. However, the comparison is based on a relatively small dataset. Lin et al. (2024a) introduces a scaling law for object diversity and environment diversity. However, it is still unclear whether 2D or 3D observation is more suitable for the scaling law.
9.9 How to enable general manipulation with dexterous capability?
Currently, foundation models for manipulation focus mainly on simple tasks like ”Pushing, Pulling, Grasping, and Placing.” Complex tasks like ”Reorientation & Relocation, Screwing, and In-hand Manipulation” have not been widely studied. The primary difference between these complex tasks and simple tasks is that these complex tasks involve dynamic contact points, high real-time requirements, and contact-rich. The primary challenge lies in the data quality and the complexity of the training process. To address this issue, it is necessary to enhance the Skill Execution module of the framework presented in this paper, with a particular focus on the Policy. However, data and training methods for dexterous manipulation are closely tied to the challenges faced in dexterous manipulation. Therefore, this section explores how to enable general manipulation with dexterous capability. It discusses the basic principles and the challenges faced of the dexterous manipulation, and also suggests improvements needed in current foundation models for manipulation to enable general manipulation to achieve dexterity.
Based on this definition, the challenges of dexterous manipulation lie in the design of the end-effector, determining a sequence of contact points and forces/torques, and whole-body control. The process of determining a sequence of contact points and forces/torques can be divided into model-based approach and the learning-based approach. The model-based approach is interpretable, explicitly showing the factors to consider in dexterous manipulation. Therefore, this section explains both model-based approach and learning-based approach. The structure of this section is shown in Fig. 9.
Figure 9: The structure of this section. First, we introduce the types and challenges of end-effector design for dexterous manipulation. Next, we introduce the factors that model-based methods need to consider. These factors provide insights for designing learning-based methods. Finally, we highlight that whole-body control is also a key factor in achieving dexterous manipulation.
9.9.1 End-effector Design.

### v5

@@ 8.4 Sim-to-Real Gap Solutions
The sim-to-real problem is a widespread issue across machine learning, not limited to manipulation (Zhao et al. (2020)). The goal is to successfully transfer the policy from the simulation (source domain) to the real world (target domain). The gap in the manipulation tasks between the simulation and the real-world includes two main types: visual gap and dynamic gap. Visual gap refers to the difference between the vision information produced by the renderer and the vision information in the real world. In terms of rendering realism, BEHAVIOR-1K (Li et al. (2023a)) highlights that Omniverse offers the highest rendering performance. The dynamic gap consists of several factors. First, there is a difference between the physics engine used in simulations and real-world physics. Second, the properties of objects, including robots, contribute to the object dynamic gap. Lastly, there is a control gap in robots, such as variations in static errors caused by different PID parameters. Currently, there are three main approaches to address sim-to-real gap: system identification, domain randomization, and transfer learning (Zhao et al. (2020)).
Most of the system identification research (Kristinsson and Dumont (1992)) aims to create an accurate mathematical model of a physical system to make the simulator more realistic. However, it is impossible to accurately build models of complex environments in simulators. The primary methods for physical parameter identification include estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)), and estimation from observations using foundation models (Gao et al. (2023)). Among these, estimation from demonstrations appears more effective. Demonstrations inherently contain interaction information and can also assist policy training. However, improving the hardware performance for collecting demonstrations remains essential.
@@
8.5 Internet-Scale Video Dataset
Methods to obtain language annotations are divided into manual and automated captions. Manual captions are created by humans labeling video content. Automated captions include four types: (1) Automatic Speech Recognition (ASR), which converts audio in videos to text (Xue et al. (2022)). (2) Alt-text, which collects captions from HTML alt-text attributes in web images and videos, like descriptions, tags, and titles (Bain et al. (2021)). (3) Transfer, which starts with a set of image-caption pairs. Then, captions are matched to video clips with similar frames (Nagrani et al. (2022)). (4) Foundation Models, which use pre-trained models to get captions. For example, VLMs provide single-frame image captions, while LLMs filter out inconsistent captions across frames (Blattmann et al. (2023)). Owing to recent advancements in language annotation techniques, most widely used internet video datasets incorporate language annotations, such as InternVid (Wang et al. (2023d)), HD-VILA-100M (Xue et al. (2022)), YT-Temporal-180M (Zellers et al. (2021)), WTS-70M (Shvetsova et al. (2025)), HowTo100M (Miech et al. (2019)), WebVid-10M (Nan et al. (2024)), and VideoCC3M (Yan et al. (2022)). At the same time, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023b)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)).
@@ 8.6 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (Mandlekar et al. (2023)) and DexMimicGen (Jiang et al. (2024c)) change the positions and orientations of objects, while CACTI (Mandi et al. (2022)) adds new, artificial objects to the scene. However, the reliability of data augmentation still needs validation. For example, MimicGen (Mandlekar et al. (2023)) filters data generation attempts based on task success. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@
Following Fig. 8, LLMs can generate credible descriptions or code for task scenes. VGMs produce 3D object meshes and render textures. Nonetheless, the validity of the generated task scenes must be ultimately assessed by VLMs. For scene generation, Automated Generation Method ensures higher diversity than Real-to-Sim. The realism of simulation depends on the simulator. Omniverse provides the best rendering performance.
@@ 9.2 What kind of learning capability should a general manipulation framework possess?
9.3 What foundation models bring for general manipulation?
The emergence of foundation models can promote the progress of general manipulation. Meanwhile, for each section, we summarize the contributions of foundation models. As for Interaction, compared to traditional methods that use fixed questioning templates to eliminate instruction ambiguity, foundation models can provide the following for ambiguous instructions and corrective instructions: 1) more natural language communication, 2) multimodal perception to detect more types of ambiguity, and 3) powerful prior knowledge to better understand user intent. As for Object Affordance and Object Recognition in Pre- and Post-conditions Detection, foundation models bring several advantages. 1) They provide the perception capabilities for open-set affordance, detection, and segmentation, enabling zero-shot recognition of novel cases. 2) The powerful prior knowledge of foundation models accelerates the learning of tool affordance. 3) Foundation models assist methods in better understanding affordance and selecting more accurate poses.
As for the hierarchy of skills: 1) Foundation models can assist in processing and interpreting natural language inputs. 2) The acquisition of world knowledge and commonsense reasoning by foundation models enhances their perception and reasoning abilities. This has the potential to improve the scalability and generalizability of tasks within the skill hierarchy. As for 3D Reconstruction and 6D Pose Estimation in State. 1) Foundation models assist in reconstructing scenes with semantic information. 2) Foundation models’ powerful 2D feature extraction ability can be applied to 3D lifting, aiding in the extraction of 3D features. 3) Foundation models enable open-set pose estimation.
As for policy. 1) Foundation models can help the model follow instructions better. 2) Foundation models can enhance the model’s generalization ability and assist reinforcement learning. 3) Foundation models trained on large manipulation data can transfer prior knowledge to new task, such as PI0 transferring the mistake correction ability of pre-trained datasets to new task. As for manipulation data generation, the main contributions of foundation models are in simulation data and data augmentation. 1) Foundation models can generate 3D mesh assets in a zero-shot manner. 2) Foundation models help create diverse simulation scene layouts. 3) The vast priors of foundation models can be applied to data augmentation.
@@ 9.4 How to use internet-scale video data for RFMs?
As for what information from video dataset can be used, there are six main types of information to convert from video datasets: (1) Pose, such as capturing human hand poses and retargeting them to dexterous hand poses (Shaw et al. (2023); Qin et al. (2022)). (2) Affordance, including grasp locations on objects and post-grasp waypoints (Mendonca et al. (2023)). (3) Motion information, explicitly includes keypoints trajectories of objects and human hand during actions (Xiong et al. (2021); Yuan et al. (2024); Wen et al. (2023b)) and implicitly includes using VQ-VAE (Van Den Oord et al. (2017)) to generate a codebook for latent delta action (Ye et al. (2024)). (4) Environment transition dynamic information, such as predicting hindsight images after completing the current action (Wu et al. (2023a); Cheang et al. (2024); Yang et al. (2023b)). (5) Semantic information, such as descriptions of current task steps (Wang et al. (2024a)) and task instruction (Jain et al. (2024)). (6) Spatial and texture information, such as MVP (Radosavovic et al. (2023)) suggests using masked autoencoding (He et al. (2022)) for improving visual reconstruction.
@@
As for how to extract these useful information, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023b)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)). When adding various labels to the video dataset, different training objectives can be used to extract features from the video dataset, such as MAE (Radosavovic et al. (2023)), contrastive learning (Ma et al. (2022)), time contrastive learning (Ma et al. (2023a)), temporal-difference learning (Bhateja et al. (2023)), video prediction objective(Du et al. (2024)), affordance prediction objective(Mendonca et al. (2023)), video-language alignment objective (Nair et al. (2022)), action motion objective(Yuan et al. (2024)) or combinations of these objectives (Karamcheti et al. (2023); Zhou et al. (2021)).
@@
As for how to utilize the extracted information to enhance or train robotic foundation models, the current robotic foundation models primarily use two learning methods: imitation learning and reinforcement learning. Therefore, the discussion on the third issue focuses on leveraging prior knowledge from video datasets in these two methods. As for imitation learning, when the robotic foundation model outputs pose and the video dataset annoated with pose label, the video dataset can be directly used as training data for the robotic foundation model (Shaw et al. (2023); Qin et al. (2022); Kareer et al. (2024)). When leveraging affordance information, motion information, environment transition dynamics information, semantic information, spatial and texture information, it is essential to employ GMM & CEM (Mendonca et al. (2023)), Inverse Dynamic Model (IDM) (Du et al. (2024); Ye et al. (2024); Wen et al. (2023b)), and Decoder (Wang et al. (2023a); Xiao et al. (2022b); Cheang et al. (2024); Wu et al. (2023a)) to transform these information into actions. Compared to other types of information, using semantic information treats the video as task instruction rather than observation (Jain et al. (2024); Shah et al. (2023); Jang et al. (2022)). At the same time, semantic information can also be used to organize tasks into a hierarchy of skills (Wang et al. (2024a)).
@@
9.5 How to uses foundation models for post-conditions detection and post-hoc correction?
The current data collection mostly focuses on gathering successful task execution data, ignoring the collection of data related to failed task executions. However, if data on failed task executions are collected and annotated with corresponding error reasons, it would be possible to train a model to both determine task execution success and analyze the reasons for task execution failure. AHA (Duan et al. (2024)) trains a VLM to assess failures and output the reasons for these failures. However, the categories of failure modes are still limited, and it cannot output more open-ended failures, such as collaboration errors in dual-arm tasks. Many current studies use internet video data to improve the generalization of policies. Exploring the use of internet video data to enhance post-condition detection and employing multimodal perception to more accurately identify the reasons for failures is a promising direction. Post-hoc correction could then generate corrective action sequences based on the reasons for task execution failure and the task objectives, which would be handed over to a policy to generate corresponding corrective actions.
9.6 How to use foundation models for End-effector Design?